We learned this in class, but I haven't yet grasped the concept. The sheet says "Solve the equation. Identify any double roots." The problem is
x^2+2=11.
I believe you're supposed to equal it to zero. So far, I have:
x^2-9=0
Then, I added zero to the other side.
X^2=9
I managed to get x=3, but I'm not positive that it's right, and I don't know what a "double root" is. Is this a double root? Thank you!

Answer by Alan3354(69443)

We learned this in class, but I haven't yet grasped the concept. The sheet says "Solve the equation. Identify any double roots." The problem is
x^2+2=11.
I believe you're supposed to equal it to zero. So far, I have:
x^2-9=0
Then, I added zero to the other side.
X^2=9
I managed to get x=3, but I'm not positive that it's right, and I don't know what a "double root" is. Is this a double root?
===============
x^2 = 9
x = 3
x = -3
----------
Not a double root, 2 separate roots.
==================
x^2 - 2x + 1 = 0
(x-1)^2 = 0
x = 1
x = 1
--------------
2 solutions, both +1.
That's a double root.